Note on ideals in polyadic numbers Milan Paštéka, Bratislava, Slovakia

Keywords: polyadic number, ideal, factor ring AMS Subject classification: 11R44 Abstract: In this paper we construct an example of an ideal in the ring of polyadic number which is not principal, has measure 0, but it is dense. We also prove that every ideal with positive measure is principal.